# mode=buildnew

### v2

@@ 3.2.1 Optimal control formulation
Analogous to how optimal control Lewis et al. (2012) can be formulated as maximum likelihood estimation in a probabilistic graphical model (Levine, 2018; Sun and Bischl, 2019), the estimation of $\theta^{(B)}$ based on the profile likelihood like objective in Equation 12 for Figure 2 is also a sequential decision process. From a statistical inference point of view, we are conducting a sequential inference (i.e., finding a sequence of $\mu$ in Equation 13), such that the last step $B$ has the best possible hyper-volume with respect to output $O$ (in a stochastic sense, see Figure 3), where $O$ can be a performance metric like accuracy on the validation set in Bayesian optimization, or defined in Equation 15 for our method.
To reformulate the estimation problem as an optimal control problem, we present the sequential decision on choosing a sequence of multipliers $\mu^{(1)},\cdots,\mu^{(B)}$ defined in Equation 1 with respect to the objective function in Equation 13 at iteration $B$ which promotes Equation 12.
$\min_{\mu^{(1)},\cdots,\mu^{(k)},\cdots,\mu^{(B)}}$ $e\mathcal{HV}_{O(\cdot)}(\theta_{\mu^{(B)}},\theta^{(0)},f_{\theta},\cdot)$ (13)
@@
$O^{(k)}=\left[\ell(\theta^{(k)}|\cdot),R(\theta^{(k)}|\cdot)\right]$ (15)
$k=0,1,\cdots,B$ (16)
Here, we use $\theta$ to denote the neural network weights, which can be regarded as the state of a dynamic system, where the state transition is governed by a low level model parameter optimization dynamic $f_{\theta}$ in Equation 14 (e.g. ADAM  (Kingma and Ba, 2015)). From a control theory point of view, $f_{\theta}$ depicts the uncertain dynamic of a plant to be controlled, while the task here is to define a controller to generate a $\mu$ sequence to guide the uncertain dynamic of $\theta$ with respect to plant output $O$, as depicted in Figure 1.We treat the loss terms $R(\cdot)$ and $\ell(\cdot)$ as the output of the system in Equation 15. From a statistical point of view, $f_{\theta}$ corresponds to the parent structure of $\theta^{(k)}$ in Figure 2, which drives $\theta$ to the next value.
Although we have no complete knowledge of $f_{\theta}$ from Equation 14 describing low level optimization algorithms like ADAM, in practice, the penalized loss from Equation 1 often descent after some number of iterations (possibly with oscillations in between), which leads to Definition 3.2.
Definition 3.2.
Therefore, we take the frequent occurrences of operators in Definition 3.2 during neural network training as a mild assumption in our following discussion. For instance, we could run $f_{\theta}$ in Equation 14 in one step to have $\ell(\theta^{(k+1)})+\mu^{(k+1)}R(\theta^{(k+1)})<\ell(\theta^{(k)})+\mu R(\theta^{(k)})$, or in more than one steps to achieve Definition 3.2.
Our final goal, however, is to ensure a joint descent of $\ell(\cdot)$ and $R(\cdot)$, which leads to Definition 3.3:
Definition 3.3.
$\ell(\theta^{+})$ $<\ell(\theta)$ (17)
$R(\theta^{+})$ $\preceq R(\theta)$ (18)
A concrete form of Equation 11 will detail how $\mu$ is updated by a controller $f_{\mu}$. Combining Equations 11 and 14, we have
@@ 3.2.2 Multi-objective optimization via shrinking the reference signal in constrained optimization
The question then arises: how can we design such a closed-loop dynamic system to ensure a Pareto descent of $O(\cdot)$ (both $R(\cdot)$ and $\ell(\cdot)$), leading to multi-objective optimization in Definition 3.3 or Equation 13?
To resolve this, we reduce it to a simpler problem, i.e., when bounding the regularization term $R(\cdot)$ from above by a time-varying reference $b^{(k)}$ (a.k.a. setpoint) in eq. 26, what value can we achieve for $\ell(\cdot)$ at Equation 25. If we could ensure the bound changes monotonically, multi-objective optimization can be achieved.
First, we define the time-varying reference signal being governed by the following:
$b^{(0)}$ $=\rho R^{(0)}$ (21)
$0<$ $\rho\in\mathbb{R}<1$ (22)
@@
where $g_{b}({\cdot})$ represents the mapping from the previous value at the training step $k-1$ to the new value at step $k$, defined as
$b^{(k)}=$ $R^{(k)},\leavevmode\nobreak\ \text{if }R^{(k)}\prec b^{(k-1)}\land\ell^{(k)}<\min_{j}\ell^{(j)}$ (24a)
$b^{(k)}=$ $j=0,\cdots,k-1$ (24b)
$b^{(k)}=$ $b^{(k-1)},\leavevmode\nobreak\ \text{otherwise}$ (24c)
@@
Note that although $b^{(k)}$ is only defined for $R^{(k)}(\cdot)$, the shrinkage of $b^{(k)}$ also depends on $\ell(\cdot)$ decrease, thus indicates multi-objective Pareto-descent in Definition 3.3.
With $b^{(k)}$ defined above, under $s_{k}$ iterations (number of iterations $s_{k}$ depends on $\theta^{(k)}$), with $b^{(k)}$ fixed, i.e. $b^{(k)}=b^{(k+1)}=\cdots=b^{(k+s_{k}-1)}$, suppose the following constrained optimization Bertsekas (2014) problem starting with $\theta^{(k)}$ has a solution and could be approximated:
$\ell^{(k+s_{k})}(\cdot)=\min_{\mu^{(k+1)},\cdots,\mu^{(k+s_{k})}}\leavevmode\nobreak\ \ell(\cdot)$ (25)
$\quad\mathrm{s.t.}\leavevmode\nobreak$ $R^{(k+s_{k})}(\cdot)\prec b^{(k)}$ (26)
$R^{(j)}(\cdot)\nprec b^{(k)}$ (27)
@@
$j=k,k+1,\cdots,k+s_{k}-1$ (29)
Then, regardless if the minimization in Equation 25 is attainable or not, as long as we have
$\ell^{(k+m)}(\cdot)$ $<\ell^{(k)}(\cdot),$ (30)
$\exists\leavevmode\nobreak\ m$ $\geq s_{k}$ (31)
, due to Equation 26, Equation 24a, we still achieve multi-objective decrease in Definition 3.3 at step $k+m$ compared to step $k$.
The evolution process of the model parameters in this case can be regarded as a plant with uncertain dynamic $f_{\theta}$ in Equations 14 and 28, where the only control we have at hand is the sequence $\mu^{(k+1)},\cdots,\mu^{(k+s_{k})}$, which provides different loss landscapes at different iterations with different $\theta$ values as depicted in Figure 4. How can we design such $\mu$ or loss landscape sequence to promote Equation 26 to happen?
To simplify the discussion, first consider $\mu,R(\cdot)\in\mathbb{R}_{+}^{d=1}$. Before the constraint in LABEL:{eq:feasibility_r} is satisfied, design $\mu^{(k+1)},\cdots,\mu^{(k+s_{k})}$ to be an increasing sequence Bazaraa et al. (2013), which gives more weight to the corresponding regularization term in $R(\cdot)$, to approximate the constraint optimization in Equation 25 by optimizing Equation 1 for each $\mu$ until Equation 26 is satisfied. However, will the increased weight on the regularization term in $R(\cdot)$ result in deteriorated $\ell(\cdot)$?
Figure 4: Illustration of the concept of multiplier induced loss landscape scheduling we proposed, used as control signal in Figure 1: The lifted landscape $\ell(\cdot)+\mu^{(k+1)}R(\cdot)$ (dotted curve) scheduled at iteration $k+1$, enables the model parameter dynamic to overcome the local minimum $\bar{\theta^{(k)}_{\mu^{(k)}}}$ of the old loss landscape $\ell(\cdot)+\mu^{(k)}R(\cdot)$ (solid curve) scheduled at iteration $k$, as can be imagined via the two balls with different colors rolling along the corresponding loss landscapes. It showcases a scenario corresponding to Definition 3.4: In comparison to $\theta^{(k)}$, $\theta^{(k+1)}$ corresponds to a decreased $\ell(\cdot)+\mu^{(k+1)}R(\cdot)$ value but increased $\ell(\cdot)+\mu^{(k)}R(\cdot)$ value.
One possibility can be, although $f_{\theta}(\mu^{(k+1)},\cdot)$ brings $\theta^{(k)}$ to the next value $\theta^{(k+1)}$ which decreases the penalized loss in Equation 1 with the new $\mu^{(k+1)}$, but the new $\theta^{(k+1)}$ evaluates to a deteriorated penalized loss with respect to the old $\mu^{(k)}$ as shown in Figure 4. To describe this behavior during the increasing process of $\mu$, we have Equation 33 in Definition 3.4. This situation is of particular interest due to the possibility of overcoming local minima of a fixed loss landscape, as indicated in Figure 4.
@@
In this situation, take $R(\cdot)\in\mathbb{R}^{d=1}$, we have the following conclusion in Corollary 3.1 where we show the increase of $\ell(\cdot)$ as an expense to decrease of $R(\cdot)$ is bounded.
Corollary 3.1.
@@
from definition 3.4 then,

### v3

@@ 3.2.1 Optimal control formulation
Continuing from Remark 3.5, to reformulate the estimation problem as an optimal control problem, we present the sequential decision on choosing multipliers $\mu^{(1)},\ldots,\mu^{(B)}$ defined in Equation 1 with respect to the objective function in Equation 6 at iteration $B$ which promotes optimization of Equation 5.
$\min_{\mu^{(1)},\ldots,\mu^{(k)},\ldots,\mu^{(B)}}$ $e\mathcal{HV}_{O(\cdot)}(\theta_{\mu^{(B)}},\theta^{(0)},f_{\theta},\cdot)$ (6)
@@
$O^{(k)}$ $=\left[\ell(\theta^{(k)}|\ldots),R(\theta^{(k)}|\ldots)\right]$ (8)
$k$ $=0,1,\ldots,B$ (9)
Here, we use $\theta$ to denote the neural network weights, which can be regarded as the state of a dynamic system, where the state transition is governed by a low level model parameter optimization dynamic $f_{\theta}$ in Equation 7 (e.g. Kingma and Ba (2015)). From a control theory point of view, $f_{\theta}$ depicts the uncertain dynamic of a plant to be controlled, while the task here is to define a controller to generate a $\mu$ sequence to guide the uncertain dynamic of $\theta$ with respect to plant output $O$, as depicted in Figure 1. We treat the loss terms $R(\cdot)$ and $\ell(\cdot)$ as the output of the system in Equation 8. From a statistical point of view, $f_{\theta}$ corresponds to the parent structure of $\theta^{(k)}$ in Figure 2, which drives $\theta$ to the next value.
Remark 3.6 (closed loop dynamic of model parameter).
A concrete form of Equation 4 will detail how $\mu$ is updated by a controller $f_{\mu}$. Combining Equations 4 and 7, we have
@@
Although we have no complete knowledge of the behavior of the low level optimization algorithms described by $f_{\theta}$ in Equation 7, the penalized loss from Equation 1 often descends after some iterations (possibly with oscillations). This leads to Definition 3.2.
Definition 3.2 (Penalized loss descent operator).
Remark 3.7 (Penalized Loss Descent Assumption).
We take the frequent existence and occurrences of such operators defined in Definition 3.2 during neural network training as a mild assumption in our following discussion. For instance, we could run $f_{\theta}$ in Equation 7 in one step to have $\ell(\theta^{(k+1)})+\mu^{(k+1)}R(\theta^{(k+1)})<\ell(\theta^{(k)})+\mu R(\theta^{(k)})$, or in more than one steps to achieve Definition 3.2.
Our final goal, however, is to ensure a joint descent of $\ell(\cdot)$ and $R(\cdot)$, which leads to Definition 3.3.
Definition 3.3.
$\ell(\theta^{+})<\ell(\theta)\quad\text{and}\quad R(\theta^{+})\preceq R(\theta)$ (12)
How can we design a closed-loop dynamic system in Remark 3.6 to ensure a Pareto descent of $O(\cdot)$, leading to multi-objective optimization in Definition 3.3 or Equation 6? To resolve this, we reduce it to a simpler problem, i.e., when bounding the regularization term $R(\cdot)$ by a time-varying reference bound $b^{(k)}$ (a.k.a. setpoint) in Equation 18, what value can we achieve for $\ell(\cdot)$ at Equation 16. If we could ensure the reference bound $b^{(k)}$ changes monotonically, multi-objective optimization can be achieved.
Thus, we define the multi-dimensional reference bound $b^{(k)}$ (setpoint) in Equations 13 to 14:
$b^{(0)}$ $=\rho R^{(0)},0<\rho\in\mathbb{R}<1$ (13)
@@ 3.2.2 Multi-objective optimization via shrinking the reference signal in constrained optimization
where $g_{b}({\cdot})$ represents the mapping from the previous value at the training step $k-1$ to the new value at step $k$, defined as Equation 15a.
$b^{(k)}=$ $R^{(k)},\leavevmode\nobreak\ \text{if }R^{(k)}\prec b^{(k-1)}\leavevmode\nobreak\ \textbf{and}$
$b^{(k)}=$ $\ell^{(k)}<\min_{j=0,\cdots,k-1}\ell^{(j)},$ (15a)
$b^{(k)}=$ $b^{(k-1)},\leavevmode\nobreak\ \text{otherwise}$ (15b)
@@
The shrinkage of $b^{(k)}$ in Equation 15a also depends on $\ell(\cdot)$ decrease, thus indicates multi-objective Pareto-descent in Definition 3.3.
Proposition 3.1 (Pareto-descent via constrained optimization with shrinking bound).
With $b^{(k)}$ defined in Equation 15a, suppose the following constrained optimization Bertsekas (2014) problem in Equations 16 to 18 starting with $\theta^{(k)}$, under $s_{k}+m_{k}$ number of iterations ($s_{k}>0$ and $m_{k}\geq 0$) has a solution
$\min_{\mu^{(k+1)},\ldots,\mu^{(k+s_{k}+m_{k})}}\leavevmode\nobreak\ \ell^{(k+s_{k}+m_{k})}(\cdot)$ (16)
$\quad\mathrm{s.t.\leavevmode\nobreak\ }R^{(j_{1})}(\cdot)$ $\nprec b^{(k)},j_{1}=k,\ldots,k+s_{k}-1$ (17)
@@
$R^{(k+j_{2})}(\cdot)$ $\prec b^{(k)},j_{2}=s_{k},\ldots,s_{k}+m_{k}$ (18)
we achieve multi-objective descent in Definition 3.3 at step $k+s_{k}+m_{k}$ compared to step $k$.
Proof.
See Appendix A. ∎
Corollary 3.1 (Approximation of constraint optimization).
Regardless of the attainability of the minimization in Equation 16, as long as we have
$\ell^{(k+s_{k}+m_{k})}(\cdot)<\ell^{(k)}(\cdot)$ (19)
we achieve multi-objective descent at step $k+s_{k}+m_{k}$.
Proof.
See Appendix A. ∎
We discuss how to approximate the constrained optimization in Section 3.2.3.
3.2.3 Loss landscape scheduling
The evolution of the model parameters can be regarded as a plant with uncertain dynamic $f_{\theta}$ in Equation 7, where the only control we have at hand is the sequence $\mu^{(k)}$, which provides different loss landscapes at different iterations as depicted in Figure 4. How can we design such loss landscape sequence to promote Equation 18 to happen? To simplify the discussion, first consider $\mu,R(\cdot)\in\mathbb{R}_{+}^{d=1}$. Before the constraint in LABEL:{eq:feasibility_r} is satisfied, design $\mu^{(k+1)},\ldots,\mu^{(k+s_{k})}$ to be an increasing sequence Bazaraa et al. (2013), which gives more weight to the corresponding regularization term in $R(\cdot)$, to approximate the constraint optimization in Equation 16 by optimizing Equation 1 for each $\mu$ until Equation 18 is satisfied. However, will the increased weight on the regularization term in $R(\cdot)$ result in deteriorated $\ell(\cdot)$?
Figure 4: Illustration of multiplier induced loss landscape scheduling, used as control signal in Figure 1: The lifted landscape $\ell(\cdot)+\mu^{(k+1)}R(\cdot)$ (dotted curve) scheduled at iteration $k+1$, enables the model parameter dynamic to overcome the local minimum $\bar{\theta}^{(k)}_{\mu^{(k)}}$ of the old loss landscape $\ell(\cdot)+\mu^{(k)}R(\cdot)$ (solid curve) scheduled at iteration $k$, as can be imagined via the two balls with different colors rolling along the corresponding loss landscapes. This showcases a scenario corresponding to Definition 3.4: In comparison to $\theta^{(k)}$, $\theta^{(k+1)}$ corresponds to a decreased $\ell(\cdot)+\mu^{(k+1)}R(\cdot)$ value but increased $\ell(\cdot)+\mu^{(k)}R(\cdot)$ value.
One possibility can be, although $f_{\theta}(\mu^{(k+1)},\theta^{(k)},\cdot)$ in Equation 7 brings $\theta^{(k)}$ to the next value $\theta^{(k+1)}$ which decreases the penalized loss in Equation 1 with $\mu^{(k+1)}$, $\theta^{(k+1)}$ evaluates to a deteriorated penalized loss with respect to the old $\mu^{(k)}$ as shown in Figure 4. To describe this behavior during the increasing process of $\mu$, we have Equation 21 in Definition 3.4. This situation is of particular interest due to the possibility of overcoming local minima of a fixed loss landscape, as indicated in Figure 4.
@@ 3.2.3 Loss landscape scheduling
3.2.4 Analysis
Based on Definition 3.4, we have the following conclusion in Proposition 3.2 where we show the increase of $\ell(\cdot)$ as an expense to decrease of $R(\cdot)$ is bounded.
Proposition 3.2 (Bounded Pareto Trade-off).
@@ 3.2.4 Analysis
from definition 3.4 assume $R(\cdot)\in\mathbb{R}^{d=1}$, then
@@
Proof.
See Appendix A. ∎
Remark 3.9 (Multi-dimensional $R$).
In the case of $R(\cdot)\in\mathbb{R}^{d>1}$, when all components of $\mu$ increase, Equation 20 and Equation 21 imply at least one component of $R(\cdot)$ will decrease. We defer to Remark 3.11 and Section 3.2.5 for the discussion on when some components of $\mu$ increase while other components decrease.
Remark 3.10 (Decrease of $\ell(\cdot)$).
Proposition 3.2 give a condition of decreasing $R(\cdot)$ at the expense of increasing $\ell(\cdot)$ (bounded though by Equation 26) at iteration $k$. When the gradient component from $\ell(\cdot)$ and $R(\cdot)$ agrees with each other at another iteration, we could expect both objectives to decrease as in Definition 3.3. It can also be the case at a particular iteration, $\ell(\cdot)$ decreases at the expense of $R(\cdot)$ increases.
What is left unclear, however, is after the accumulative effect of several iterations, whether these exists an $s_{k},m_{k}$ ensuring Equation 19. To answer this, we derive Proposition 3.3, which gives a bound on the accumulative change of $\ell(\cdot)$ with $B=s_{k}+m_{k}$.
Proposition 3.3 (Single-step $\ell$ increase bound).
Suppose that for all $k=0,\dots,B-1$
and $\mu^{(k+1)}>\mu^{(k)}$. Then
$\ell(\theta^{(B)})\leq\ell(\theta^{(0)})+S_{>}+S_{<},$ (27)
where
$S_{<}$ $=\sum_{k\in\mathcal{K}_{<}}\mu^{(k+1)}(R(\theta^{(k)})-R(\theta^{(k+1)})),$
[…]
$\mathcal{K}_{<}$ =$=\{0,\dots,B-1\}\setminus\mathcal{K}_{>}.$
Proof.
See Appendix A. ∎
Based on Proposition 3.3, we have the following conjecture to ensure Equation 19
Conjecture 3.1 (Multi-step $\ell$ decrease).
Further decompose $S_{<}$ from Proposition 3.3 into $S_{<}^{-}$ and $S_{<}^{+}$ via decomposing $\mathcal{K}_{<}$ in Proposition 3.3 into $\mathcal{K}_{<}^{-}$ and $\mathcal{K}_{<}^{+}$. Here $\mathcal{K}_{<}^{-}$ corresponds to the situations of $\ell(\cdot)$ descent with $S_{<}^{-}<0$. And $\mathcal{K}_{<}^{+}$ corresponds to the $\ell(\cdot)$ ascent cases but upperbounded by $S_{<}^{+}=\sum_{k\in\mathcal{K}_{<}^{+}}\mu^{(k+1)}(R(\theta^{(k)})-R(\theta^{(k+1)}))$. Then $\exists$ sequence $\{\mu^{(k)}\}$ s.t.
$\ell(\theta^{(B)})-\ell(\theta^{(0)})\leq S_{>}+S_{<}^{+}+S_{<}^{-}<0$ (28)
Remark 3.11 (Trade-off Multi-dimensional $\mu$).
In the case of $d>1$, as long as one loss component of ${R}^{(k)}(\cdot)$, without loss of generality, say $R_{1}^{(k)}(\cdot)$ is not bounded by its corresponding setpoint component $b_{1}^{(k)}(\cdot)$, we can still increase the value of the $\mu$ component $\mu^{(k)}_{1}$ to give more weight to the corresponding loss component, until the loss component in question decreases below the setpoint. However, an increased $\mu$ component $\mu^{(k)}_{1}$ corresponds to more weight on the gradient component corresponding to $R_{1}^{(k)}(\cdot)$, which makes it harder for other components of $R(\cdot)$ and $\ell(\cdot)$ to decrease. Therefore, instead of always increasing each component of $\mu$, if one component of $R$ has overshoot over the setpoint (i.e. constraint switched from being not satisfied to satisfied with respect to a particular loss component), we decrease the $\mu$ value to give the other components of $\mu$ and $R$ more feasible space to adjust themselves.
Remark 3.12 (Multi-step Pareto Descent).
We could extend 3.1 to $\exists\mu$ sequence such that after $B$ iterations, Pareto descent in Definition 3.3 could be achieved (equivalently, setpoint should shrink see Remark 3.8). We did observe such a Pareto descent behavior in our experiment for the $d>1$ case of Equation 40, e.g. see Figure 5 and Figure 6.
1:procedure M-Hof-Opt($\mu^{(0)},\theta^{(0)},\rho,\cdot$)
2:     Initialize $\mu^{(0)}$, calculate $b^{(0)}=\rho R^{(0)}$.
3:     Compute $K_{I}$ based on Remark 3.13.
4:     while budget $B$ not reached do
5:         Update $\mu$ from controller in Equation 31.
6:         Update $\theta$ via Equation 7 with updated $\mu$.
7:         Adapt setpoint according to Equation 14.
8:     end while
9:     return $\theta^{(B)}$triangleright
10:end procedure
3.2.5 Output feedback PI-like controller
Based on Remark 3.11, with the adaptive law from Equations 14 and 15a for setpoint $b$, we design the following controller (see Figure 1) for $\mu$:
$e^{(k)}$ $=R(\cdot)-b^{(k)}$ (29)
$\delta_{I}^{(k+1)}$ $=(1-\xi_{d})\delta_{I}^{(k)}+\xi_{d}e^{(k)}$ (30)
$K_{I}$ $>0$ (32)
$K_{I}$ $\in\mathbb{R}^{d}$ (33)
$k$ $=0,1,\ldots,B$ (34)
In Equation 29, we calculate how far the current output is away from the setpoint $b$, which gets passed through a moving average in Equation 30 with $\xi_{d}$ being the coefficient of moving average in Equation 30 (Rezende and Viola, 2018). $\mu$ is the output of the controller in Equation 31, $K_{I}$ is the control gain for PI (proportional-integration) (Johnson and Moradi, 2005) like control, $K_{I}\delta_{I}^{(k+1)}$ is component wise multiplication, $\exp$ and $\max$ are computed component wise, $v_{sat}$ is the exponential shoulder saturation, which defines the maximum rate of change of multipliers. $\mu_{clip}$ is the upper bound for $\mu$. Note that $\mu_{clip}$ determines the dynamic range of $\mu$ while $v_{sat}$ determines an upper bound about how fast $\mu$ changes.
Remark 3.13.
The whole process is summarized in Algorithm 1, which details Figure 1.
3.2.6 Multi-objective setpoint based model selection
3.3 Related work and discussion
We discuss the advantages of our method compared to other multiplier scheduling and multi-objective optimization methods in Appendix B.
4   Experiments
We conduct experiments and benchmarks to answer the following questions: Is M-HOF-Opt capable of adjusting multidimensional multipliers automatically to drive each component of $R(\cdot)$ at different scales to the setpoint? Does the setpoint shrink, which implies multi-objective descent? See Figure 5 and Figure 6 in Section 4.1.
Will wrong multiplier choices for warmup and fixed multiplier training schemes result in catastrophic effects? In contrast, does varying controller hyperparameters have detrimental effects on M-HOF-Opt? See Figure 7 in Section 4.2, as well as Figure 8 in Section C.4.
Does changing the lower-level optimization algorithm (i.e. the open-loop plant or the feedback-free model parameter dynamic system M-HOF-Opt has to control) have an impact on the performance of M-HOF-Opt? See Section C.5.
With respect to the above questions, we got favorable results for M-HOF-Opt in all experiments which we analyze below. The experimental and benchmark setting is detailed in Section C.1 with implementation in https://github.com/marrlab/DomainLab/tree/mhof from DomainLab (Sun et al., 2024).
Figure 5: Our method drives different loss terms of $R(\cdot)$ at different scales and rates towards the setpoint, which further promotes the setpoint shrinkage. In the Top row, we show the multiplier dynamic as controller output signal in Equation 31 across training epochs. In the Bottom row, we present the tracking behavior of the corresponding regularization loss $R(\cdot)$ in Equation 40 with respect to setpoint $b$ defined in Equation 14.
4.1 Illustration of multi-objective descent
Since the domain generalization model DIVA (Ilse et al., 2020) with math formulation explained in Section C.2.1 has 6 loss terms [^1] At time of writing, we are not aware of any other deep learning model with more loss terms., it is a good example to demonstrate the advantage of our algorithm in trading off many loss terms. Our hierarchical output feedback optimization training scheme produced the dynamics of $R(\cdot)$ shown in Figure 5 together with the corresponding setpoints $b$ and multipliers. Note that the multiplier $\beta_{y},\gamma_{d},\mu_{recon}$ and their corresponding loss components operate at different numerical ranges, but our controller still manages to drive the different $R(\cdot)$ loss terms down at different scales and rates.
Remark 4.1 (setpoint shrinkage).
Regarding the dynamics of how the setpoint $b$ defined in Equations 13 to 14 decreases in Figure 5: Initially, Equation 17 holds, in which case we do not update the setpoint, even if some but not all the blue curves (corresponding to $R(\cdot)$ components) are below the red curves (components of the setpoint). After all constraints for each component of $R(\cdot)$ are satisfied as described in Equation 18, the setpoint (red curves) gets adapted to a new value (which we term shrinkage), as a new goal to be reached by the $R(\cdot)$ loss in Equation 46. Note that the shrinkage of setpoint depends on Pareto dominance in Equation 15a, while non-dominance defined an equivalent relation in Definition 2.2.
Figure 6 shows the output portrait of $\ell(\cdot)$ versus a component of $R(\cdot)$ of our training scheme This confirms the multi-step Pareto descent discussed in Remark 3.12.
With the multi-objective setpoint based model selection criteria in Section 3.2.6, our method admits a multi-objective descent of the selected model compared to the initial output, such that the uncontrolled behavior at the end of iterations can be safely ignored.
Figure 6: Our training scheme adapts the multi-dimensional multiplier $\mu$, steering the optimization process towards a configuration that balances the trade-off between minimizing output $\ell(\cdot)=\mathbb{E}_{q_{\phi_{y}}(z_{y}|x)}[\log q_{\omega_{y}}(y|z_{y})]$ in Equation 41 and output $R_{\gamma_{d}}(\cdot)=\mathbb{E}_{q_{\phi_{d}}(z_{d}|x)}[\log q_{\omega_{d}}(d|z_{d})]$ in Equation 46. In this plot, the solid large point corresponds to the initial output. We use gradually changing colors to indicate the training iterations (epochs), indicated by the color bar. For improved visualization, we exclusively utilized the initial 120 epochs and plotted data points solely for every 10 epochs. This figure corresponds to the experimental setting in Figure 5.
4.2 Benchmark results
We demonstrate the power of M-HOF-Opt in automatic adaptation of multipliers via benchmarking out-of-domain generalization performance in Figure 7 for training DIVA Ilse et al. (2020) compared to baseline training schemes. From Figure 7, we can observe that the generalization performance of baselines is highly sensitive to multiplier combination. In contrast, our method, as supported by the adaptive behaviors visualized in Figure 5, automatically adjusts the multipliers during the training process, leading to a robust performance with respect to changes of controller hyperparameters and multiplier initial condition $\mu^{(0)}$. Our method does not need hyperparameter searching of the multiplier, thus saves a fair amount of computational resources. For additional experiments, see Section C.4.
5   Conclusion
This work uses control theory to address the issue of combinatorial choice for multidimensional multipliers weighting many loss terms in structural risk minimization for deep neural networks by proposing a novel multi-objective optimization algorithm via constraint optimization with shrinking bound. We develop a probabilistic graphical model for joint model parameter and multiplier optimization with respect to a multi-objective descent of all loss terms. The estimation of model parameter leads to an automatic adjustment scheme for the multipliers, adopting a hierarchical control scheme, which breaks the multi-objective optimization problem into a series of constraint optimization sub-problems. Each sub-problem is configured with a self-adaptive multi-objective setpoint updated via Pareto dominance. A PI-like multiplier controller drives the loss term to satisfy the setpoint constraint for each sub-problem. Our method operates at the timescale of epoch level during training, thus circumvents the need for exhaustive multiplier search and saves tremendous computational resources compared to methods like Bayesian optimization. It also circumvents the excessive memory requirements and heavy computational burden of existing multi-objective deep learning methods. Our method demonstrates robust out-of-domain generalization performance against controller hyperparameter variation compared to other multiplier choice or scheduling schemes which produce very unstable behavior in the training due to the need for a combinatorial choice of multipliers.
Acknowledgements
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes: See Section 2, Section 3Section C.2 ]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes: See Appendix B ]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Not Applicable to camera-ready submission]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes: See Section 3 ]
    2. (b)
      Complete proofs of all theoretical results. [Yes: See Appendix A ]
    3. (c)
      Clear explanations of any assumptions. [Yes: See Section 3 ]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes: See Section 4Section C.1 ]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes: See Section C.1 ]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes: See caption of Figure 7 ]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes: See Appendix D ]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes: See Section C.1Appendix E ]
    2. (b)
      The license information of the assets, if applicable. [Yes: See Appendix E ]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes: See Section 4 ]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Appendix
Appendix A Mathematical proofs
In this seciton, we restate theoretical results in the main text with the same statement index and provide proofs. See 3.1
Proof.
Here $s_{k}>0$ is the number of iterations maintaining constraint in Equation 17, and $s_{k}+m_{k}$ is the number of iterations to ensure the decrease of $\ell$ subject to decreasing $R$ in Equation 18.
Note that $b^{(k+j)}$ remains fixed before $s_{k}$, i.e. $b^{(k)}=\cdots=b^{(k+s_{k}+m_{k}-1)}$ due to Equation 15a. Between $s_{k}$ till $s_{k}+m_{k}$, due to Equation 15a, $R$ has decreased.
Due to Equation 18, the new reference bound $b^{(k+s_{k})}$ (setpoint) has been respected, so $R^{(k+s_{k}+j)}$ with $0\leq j\leq m_{k}$ still remains better than $R^{(k)}$, even though it can be worse than intermediate values $R^{(k+s_{k}+j^{{}^{\prime}})}$ with $0<j^{{}^{\prime}}<j$.
Now due to Equation 15a, the minimization of $\ell^{(k+s_{k}+m_{k})}$ cannot lead to values worse than $\ell^{(k)}$ because otherwise Equation 18 cannot be respected.
Thus both $\ell$ and $R$ decrease at step $k+s_{k}+m_{k}$ compared to step $k$.
See 3.1
Proof.
Feasibility in Equation 19 and the condition in Equation 18 ensures Pareto-descent in Definition 3.3, following the proof in Proposition 3.1. ∎
See 3.2